# Dynamical phases of short-term memory mechanisms in RNNs

### v1

@@ 3.2 Influence of task design on the learned latent mechanism
Behavioral experiments in systems neuroscience often contain arbitrary design choices that are deeply embedded in the experimental paradigm, yet their implications are rarely scrutinized. Consider the delayed cue discrimination task, which is similar to the delayed activation task, but with two input and output channels where the delay period begins after a brief cue presentation window ($T_{\rm in}$, discussed below in Fig. 6). While animals typically indicate their response by licking either a right or left spout, seemingly minor procedural decisions – such as whether to provide immediate feedback or require a short post-response waiting period – can fundamentally alter the mechanisms networks learn to solve the task.
Our investigations reveal that implementing a post-reaction period qualitatively changes the dynamical mechanisms emerging in trained networks. In our delayed activation task, we already assume that a previous cue has put the network in a stereotypical state, hence the initialization to a particular state in the task. Then, we can consider our toy tasks as simplified versions of cue discrimination tasks, where we focus on the retention of a particular cue. With this in mind, in networks trained without a post-reaction period ($T_{\rm post}=0ms$), the dynamics typically converge to stable point attractors, creating a straightforward mapping between stimuli and responses. However, when we introduce even a brief post-reaction period ($T_{\rm post}=30ms$), the networks predominantly learn limit cycle solutions, despite achieving similar task performance (Fig. 4A). The phase portraits in Fig. 4B clearly illustrate this distinction, showing fixed point convergence in networks without a post-period versus circular trajectories indicating limit cycles in networks with a post-period.
[…]
This finding becomes particularly intriguing in the context of the current emphasis on dynamical systems approaches in neuroscience, where considerable attention has been given to identifying and manipulating ([39, 40]) attractors and manifolds underlying neural activity. The stark difference in dynamical solutions – stable points versus limit cycles – emerging from such a minor task modification raises important questions about the interpretation of dynamical structures observed in neural recordings. Can we definitively claim that one dynamical solution is more fundamental than another simply because we observed it in animals performing a particular variant of a task?
To systematically investigate this phenomenon, we conducted a broader empirical study examining how delay period duration interacts with the presence of a post-reaction window (Fig. 4C). Our results demonstrate that increasing the delay period generally promotes the emergence of limit cycle solutions, but this tendency is dramatically amplified by the addition of a post-reaction window. Statistical analysis using Fisher’s exact test with Bonferroni corrections confirms the significance of these effects across different delay periods. These findings serve as a crucial reminder about the profound influence of experimental design choices on the neural mechanisms we observe and interpret. They suggest that we must exercise greater caution when drawing conclusions about the fundamental nature of neural computations based on specific task implementations, as the observed dynamical structures may be more contingent on subtle task parameters than previously appreciated.
3.3 A theoretical scaling analysis of mechanistic phases via toy dynamical system models
@@
So far, we used simple examples to demonstrate how RNNs can use different latent mechanisms to generate sequences, that the choice between mechanisms is not random, and that RNNs can change their inner mechanisms during the course of their training. Now, we turn to a set of toy models and provide theoretical insights into why these phenomena occur.
@@
Previous work has considered a simple toy dynamical system model, the normal form for the saddle-node bifurcation [41], given as $\dot{x}=x^{2}+r$ for a trainable $r\in\mathbb{R}$ parameter [37]. Earlier work studied the complex learning dynamics in a simple system that can generate slow-points. In this section, we set out to extend this analysis to a second toy model, a limit cycle attractor. But first, in Appendix S1.1, we reproduce the findings of the earlier work for a more general version of the delayed-activation task with $T_{\rm delay}\neq T_{\rm resp}$. This analysis reveals a maximum learning rate, also reported in [37], beyond which the learning dynamics destabilize and even an approximation of the desirable solution cannot be reached. Our analysis reveals that the learning rate scales following a power law such that $\alpha_{\rm SP}\sim O(T_{\rm delay}^{-\beta})$, where $\beta\in[4,5]$. Such a steep decrease with the desirable learning rate can explain our anecdotal observations in Fig. 2 that slow-points become rarer for large $T_{\rm delay}$ or high learning rate values.
Now, we introduce a new toy model that models the limit cycles that follows a two-dimensional dynamical system with periodic orbits. For simplicity, we fix the radius to an attractive fixed-point, and allow a periodic oscillation in the frequency such that the equations of the motion for the toy model become (See Appendix S1.2 for derivations):
$x(t)=\sin(2\pi rt),$ (3)
@@
where $r\in\mathbb{R}$ is a learnable parameter and the network output is defined as $\hat{o}(x)=\Theta(x<0)$, where $\Theta(\cdot)$ is defined as the Heaviside function. By inspection, the rotations have a half-period of $T_{\rm half}=\frac{1}{2r}$, and the model outputs zero for the first $T_{\rm half}$ while outputting one for the second half. An example flow map for this model is shown in Fig. 5A. Changing the parameter $r$ simply updates the speed of the limit cycle, though as we show below, this is not a trivial task.
Figure 5: A theoretical scaling analysis for algorithmic phases. A The illustration of the flow map corresponding to the toy limit cycle model for $r=0.1$. The red circle denotes the attractive limit cycle, the purple line corresponds to the decision boundary of the output, and the green circle denotes the starting point of the state. B An example loss curve for the limit cycle toy model, with analytical curve and empirical data points. Note the kink at the global minimum. C The different dependencies of the effective learning rate for two types of mechanisms creates a phase space of mechanisms as a function of delay and learning rate values. For this figure, we picked $T_{\rm resp}=5$. In general, there are three main phases of algorithmic mechanisms for large $T_{\rm delay}\gg T_{\rm resp}$.
@@
In Appendix S1.2, we compute an analytical loss function for the limit cycle model in the limits $T_{\rm resp}=T_{\rm delay}$ and $T_{\rm resp}\ll T_{\rm delay}$. We illustrate the latter case in Fig. 5B, note the exact correspondence between the analytical curve and the numerical simulations. For both cases, the loss function has a global optimum at $r=\frac{1}{2T_{\rm delay}}$. However, it has a kink around the global minimum and quickly flattens to a constant value for $r\leq\frac{1}{2(T_{\rm delay}+T_{\rm resp})}$. Hence, if the learning rate is larger than some critical value, the oscillations around the global minimum can throw the parameter $r$ to a value $r<\frac{1}{2(T_{\rm delay}+T_{\rm resp})}$, terminating the training process due to the zero gradient despite the fact that the model cannot solve the task. Our analysis reveals that this learning rate scales following $\alpha_{\rm LC}^{*}\sim T_{\rm delay}^{-\gamma}$ where $\gamma\in[2,3]$. Notably, the limit cycle solution has a better scaling than the slow-point one, providing a theoretical explanation for our observation in Fig. 3. Specifically, since the RNN in Fig. 3 was trained using a learning rate right at the transition (Compare to Fig. 2), it is not surprising that a slow point mechanism eventually evolves into a limit cycle. Moreover, this theoretical prediction is also consistent with the experiments in Fig. 4C, where increasing the delay period for a fixed learning rate resulted in more RNNs finding limit cycle based solutions. Similarly, by simply shifting the decision boundary from $x<0$ to $x<-x^{*}$ for some $x^{*}\leq 1$, this model can also incorporate a post-reaction period of zero outputs, for instance by setting the total period equal to $T_{\rm delay}+T_{\rm rep}+T_{\rm post}$.
At least as interesting is the fact that the different scaling of these solutions defines a phase space of mechanisms that can be learned by RNNs. This phase space, illustrated in Fig. 5C, depends on the delay time (a task parameter) and the learning rate (an optimization parameter). Below, we show empirically that similar phase diagrams emerge in full-rank RNNs trained on delayed cue discrimination tasks. It is worth noting that increasing the dimensionality of the dynamical system can allow more efficient solutions to the system, but the toy models we discuss can be thought as approximate bounds on what can be achieved.
3.4 A large-scale analysis with RNNs reveals an empirical scaling law and mechanistic phases
Figure 6: Large scale experiments with over 35,000 RNNs performing the delayed cue discrimination tasks. A-B An illustration of the delayed cue-discrimination task and simulation plots. The task includes two correct cue-reaction matches (blue-left, red-right) and consists of four periods: (i) Cue period, where one of the two cues is presented; (ii) Delay period, with no expected reaction; (iii) Reaction period, where the cue class is expected; and (iv) an optional Post-reaction period at the end of the trial. C-D The distribution of different learning rates and the resulting memory mechanisms depending on the delay period. Adding a post-reaction period restricts the model to learn limit cycles. Experiments start with different seeds. E-F The plots illustrate the fitting of the power law in a log-log scale, in which the y-axis corresponds to the maximum learnable delay and x-axis to the critical learning rates. Each dot represents the cutoff delay value for a particular learning rate, indicating the point beyond which learning ceases (See Methods for how we estimate these delay values). As predicted, both scenarios illustrate scaling laws between the learnable delay values and the learning rates. While the original delayed cue discrimination task provides a critical exponent within the theoretically predicted values $\beta^{*}\approx-2.40$, the modified version allows more efficient training with $\beta^{*}\approx-1.57$. Parameters: $T_{\rm in}=30ms$, $T_{\rm resp}=50ms$, $\tau=10ms$, $\Delta t=5ms$, iid noise with $10^{-3}$ s.d., the delay window, the learning rates, and post-reaction times were varied. For panels C-D, we trained slightly more than 25,000 RNNs, whereas for panels D-F, we had roughly 5,000 RNNs.
Figure 7: Emergent mechanistic phases as a function of delay and learning rates. We plotted the results of the experiments in Fig. 6C-D in the form of a phase diagram, where the colors were proportionally scaled to match the three possible outcomes. As predicted by the theoretical analysis in Fig. 5C, three phases can be clearly observed in the original task with no post-reaction window (left), whereas adding the post-reaction window leads to limit cycle solutions (right).
Next, we tested our theoretical predictions by training over 35,000 RNNs on delayed cue-discrimination tasks (Figs. 6 and S1). Briefly, delayed cue-discrimination task consists of four distinct periods: cue, delay, reaction, and an optional post-reaction period (i-iv). Upon observation of a cue, the networks have to wait for a brief period and output a response in the corresponding output channel. If there is a post-reaction period, then the networks should learn to output zeros after the response window concludes.
First, we examined whether different mechanisms would consistently emerge in different values of the delay windows, as predicted by our theoretical analysis earlier. In Fig. 6 (C-D), we observed that increasing the length of the delay period preferentially led to the emergence of limit cycles, consistent with our findings in Figs. 2 and 3. Yet, RNNs still developed slow-point manifolds for shorter delays, which was nonetheless absent for RNNs performing a modified version of the task with a post-reaction window (Fig. 6 (C-D)). Notably, when we trained on a version of the task that no longer required a memory maintenance component (Fig. S2), the dependence on the full length of the trial was minimal and the relationship observed in Fig. 6C-D was practically ablated. Therefore, overall, these results suggested that task requirements shape the viability of different memory strategies.
Next, we quantified the relationship between delay duration and learning performance through scaling law analysis (Fig 6E-F). The critical exponent $\beta$ differed significantly between the original task ($\beta\approx-2.40$) and the modified version with post-reaction period ($\beta\approx-1.57$). Interestingly, however, the critical exponent extracted from the former was within the expected range of values $\beta\in[2,3]$ from our theoretical studies using the delayed activation task. Finally, to comprehensively map these mechanistic transitions, we plotted a phase diagram (Fig 7) that captures both the emergence of different memory mechanisms and their associated scaling laws.
@@ 4 Discussion
Learning to respond based on delayed information is a fundamental challenge in both neuroscience and machine learning [42, 43]. In machine learning, early studies in the 1990s identified this difficulty as the problem of learning long-term dependencies, particularly in recurrent neural networks (RNNs) trained on tasks requiring information retention across time [43]. Despite significant progress in neural network architectures, this issue remains central to designing systems capable of effective temporal reasoning.
To address long-term dependencies, researchers developed specialized architectures such as Long Short-Term Memory (LSTM) networks, Gated Recurrent Units (GRUs), and, more recently, Transformers [44, 45, 46]. These models introduced mechanisms that improved memory retention by adjusting internal timescales. However, unlike biological neurons, which operate under fixed timescales—such as the “membrane time constant,” which governs the decay of neuronal firing [47, 48, 49, 50, 51]—these architectures learn their own timescales dynamically. This fundamental difference raises the question: why do biologically inspired networks struggle to learn long-term dependencies, and what constraints govern their performance?
@@
A key insight into this question comes from examining the relationship between learning rate and delay length. In simple delayed-response tasks, we show that the maximum learning rate, beyond which training fails, follows a power-law scaling with delay length. This implies that as the delay increases, learning requires increasingly small updates, ultimately rendering training impractically slow for long delays. This limitation may have played a role in capturing long-term dependencies, which was the focus of the early RNN studies [43].
The development of gated architectures such as LSTMs and GRUs helped alleviate this issue by introducing adaptive timescales, allowing networks to modulate how past information is retained [44]. In principle, flexible timescales can overcome the power-law limitation of fixed dynamical systems. However, this comes at a cost. Extending timescales indefinitely would slow down computation, making rapid decision-making difficult. Moreover, increasing the parameter space to accommodate longer memory spans complicates optimization, potentially destabilizing learning [37]. Unlike biological networks, which are constrained by predefined timescales such as action potentials, artificial networks rely on optimizing these scales through training—a difference that influences their ability to generalize across temporal tasks.
Understanding these fundamental constraints provides a new perspective on why learning long-term dependencies is difficult across different architectures. By framing the problem through the lens of dynamical systems theory, we characterized a crucial trade-off as part of this problem. The identification of this trade-off may allow future work to explore potential solutions that balance biological plausibility, computational efficiency, and learning performance.
Our work also has implications for experimental neuroscience. Specifically, limit cycles have been studied in terms of phase matching (i.e., periodic computation) [52]. But, here we have shown that more generally, they can also be used to approximate sequential neuronal activations. However, in a dynamic brain, where synaptic plasticity changes functional connectivity rapidly [36], how can we figure out whether slow points or limit cycles subserve observed neural dynamics? More broadly, how can we tell apart solutions that locally form a class of equivalence, but are globally distinct. To the best of our knowledge, this unique concept has not been addressed or raised as part of the experimental design in systems neuroscience research. Yet, addressing these questions would require a new style of experiments to be designed. To start with, in this work, we illustrated how different manipulations in the task structure can shed light into dissociating these mechanisms.
Finally, beyond neuroscience, this phenomenon of staged learning dynamics, algorithmic phase transitions, is widely observed in artificial networks including transformers trained on mod-arithmetic tasks [53, 54], language modeling [55, 56], and in-context learning [57]. In these studies of learning dynamics on sequential tasks, our study of RNNs offers a unique perspective on interpretability by connecting the computational mechanisms of circuits to foundational mechanisms in dynamical systems. Furthermore, these recent works [58] aim to identify progress measures of learning, akin to order parameters in physics, to see if phase transitions occur. Here, we see that literal phases emerge with distinct computational mechanisms. Future works could explore ways to functionally map those dynamical mechanisms of computation to mechanisms implementable by feedforward architectures like Transformers.
5 Conclusion
In this work, we explored the mechanisms of how recurrent neural networks developed and utilized different mechanisms for maintaining short-term memory. Through extensive theoretical analysis and empirical validation with over 35,000 trained RNNs, we identified two distinct memory maintenance strategies that subserved the emergence of neural sequences: slow-point manifolds and limit cycles. We demonstrated that the emergence of these mechanisms followed scaling laws relating learning rates to delay periods, and showed how subtle changes in task structure (like adding a post-reaction period) could fundamentally alter which mechanism predominated. Our findings have important implications for both machine learning and neuroscience - they help explain why learning long-term dependencies is inherently difficult in neural networks with fixed timescales, and they suggest new experimental approaches for investigating memory mechanisms in biological neural circuits. Most importantly, our work revealed that neural circuits may use multiple, mathematically equivalent mechanisms to solve the same memory task, highlighting the need to carefully consider task design when interpreting neural recordings.

### v2

@@ 3.2 Influence of task design on the learned mechanism
Behavioral experiments in systems neuroscience often contain arbitrary design choices that are deeply embedded in the experimental paradigm, yet their implications are rarely scrutinized. Consider the delayed cue discrimination task, a task commonly studied in neuroscience experiments with mice (Ebrahimi et al., 2022). In the machine learning context, it consists of four distinct periods (Fig. 1D-E): cue, delay, reaction, and an optional post-reaction period (i-iv). Upon observation of a cue, the networks have to wait for a brief period and output a response in the corresponding output channel. Typically, there are two possible cues and hence two output channels. If there is a post-reaction period, then the networks should learn to output zeros after the response window concludes. In experimental context, while animals typically indicate their response by licking either a right or left spout, the reward or punishment is often directly provided at the conclusion of the behavior without the post-reaction window. The distinction, as we show below, matters for the final learned solution.
Minor modifications to the task requirements can change the learned mechanisms–We now set out to answer our second question (i.e., Q2: What determines which mechanism is favored under different task or training conditions?). To do so, we perform a minor change to the delayed activation task by requiring a short post-reaction waiting period and illustrate how minor changes in experimental design can fundamentally alter the mechanisms networks tend to learn during training. In the modified delayed activation task, (rank-2) RNNs are required to switch their responses at the end of the response period and output zero for a brief duration before the trial concludes (Fig. 1D-E). In networks trained without this modification ($T_{\rm post}=0ms$), the latent dynamical systems can learn both slow-point manifolds (for shorter delays) and limit-cycles (for longer delays), similar to Fig. 2. However, when we introduce even a brief post-reaction period ($T_{\rm post}=30ms$), the networks predominantly learn limit cycle solutions, despite achieving similar task performance (Fig. 3A-B). We quantified and confirmed these findings with a broader empirical study with $1600$ RNNs (Fig. 3C; via Fisher’s exact test with Bonferroni corrections). Thus, in rank-2 RNNs solving the delayed activation task, increasing the delay period promoted the emergence of limit cycle solutions, and this tendency was dramatically amplified by the addition of a post-reaction window. This demonstrated that seemingly minor changes in the task design can favor learning distinct short-term memory maintenance mechanisms.
3.3 A theoretical scaling analysis of mechanistic phases
@@
So far, we used simple examples to demonstrate how RNNs can use different latent mechanisms to generate neural sequences, that RNNs can change their inner mechanisms during the course of their training, that the choice between these mechanisms is not random but depends on specifics of the experimental design. Now, to provide an answer to the third question (i.e., Q3: How do the learned mechanisms vary with delay duration?), we turn to a set of low-dimensional dynamical models and study their learning dynamics.
Why study low-dimensional toy models?–Recent work has shown that training (full-rank) RNNs often converges to effectively low-dimensional solutions (Valente et al., 2022), a property that appears to generalize across many real-world networks (Thibeault et al., 2024). As discussed in Eq. (2), these low-dimensional computations can often be interpreted as latent dynamical systems. In this view, training the RNN parameters $(W,W^{\rm in},b)$ amounts to shaping a latent flow map of the form:
$\tau\dot{\kappa}(t)=G(\kappa(t),u(t);W,W^{\rm in},b),$ (3)
where $G$ governs the evolution of low-dimensional latent variables $\kappa(t)$ in response to input $u(t)$. Hence, studying how toy models, which do not necessarily share the same architectural assumptions as RNNs but still aim to learn similar local dynamics, may provide intuition for how these flow maps may be learned. Thus, we set out to analyze simple low-dimensional dynamical systems that capture the two mechanisms (slow-points and limit cycles) we have observed in trained RNNs.
Earlier work has considered a one-dimensional toy system, the normal form of a saddle-node bifurcation, $\dot{x}=x^{2}+r$ (Strogatz, 2018), to study the learning dynamics of slow-point formation (Dinc et al., 2025b). This system approximates the local behavior of one-dimensional flow maps $G(\kappa,u)$ near their minima, where slow-points (also observed in Fig. 2) arise. To perform a similar analysis with the oscillatory dynamics observed in Figs. 2 and 3, we will consider a second toy model in this work: a two-dimensional system exhibiting a limit cycle, one of the simplest settings in which periodic orbits, and hence oscillations, can emerge (Strogatz, 2018). Both toy models are meant to qualitatively mirror the behavior of the respective latent flow maps $G(\kappa,u)$ that may be learned by an RNN (Fig. 2). While these toy systems are not exhaustive, they capture key features of the latent mechanisms that RNNs may converge to, and allow us to dissect their learning dynamics in a controlled and analytical setting, which we discuss next.
@@
Extracting scaling laws from toy models–To start with, in Appendix S2.1, we reproduce the findings of the earlier work on slow-point manifolds for a more general version of the delayed activation task with $T_{\rm delay}\neq T_{\rm resp}$. This analysis reveals a maximum learning rate, also reported in (Dinc et al., 2025b), beyond which the learning dynamics destabilize and even an approximation of the desirable solution cannot be reached. Our analysis reveals that this maximum learning rate follows a power law such that $\alpha_{\rm SP}\sim O(T_{\rm delay}^{-\beta_{\rm SP}})$, where $\beta_{\rm SP}\in[4,5]$. The exact value $\beta_{\rm SP}$ achieves within this interval depends whether $T_{\rm delay}$ and $T_{\rm resp}$ are comparable, with larger delays leading to more substantial decreases (approaching $\beta_{\rm SP}\to 5$ in the limit $T_{\rm delay}\gg T_{\rm resp}$). Such a steep decrease with the desirable learning rate can explain our anecdotal observations in Figs. 2 and 3C that slow-points become rarer for large $T_{\rm delay}$ or high learning rates, which will be substantiated in the next section.
@@
Now, we focus on the limit cycle model. For simplicity, we fix the radius to an attractive fixed-point, and allow a periodic oscillation in the frequency such that the equations of the motion for the toy model become (See Appendix S2.2 for derivations) $x(t)=\sin(2\pi rt)$, where $r\in\mathbb{R}$ is a learnable parameter. The network output is defined as $\hat{o}(x)=\Theta(x<0)$, where $\Theta(\cdot)$ is defined as the Heaviside function. By inspection, the rotations have a half-period of $T_{\rm half}=\frac{1}{2r}$, and the model outputs zero for the first $T_{\rm half}$ while outputting one for the second half. An example flow map is shown in Fig. S2A; changing the parameter, $r$, modulates the frequency/period of the cycle.
Figure 4: Emergent phases of attractor mechanisms in RNNs performing delayed cue-discrimination. A Loss curves corresponding to different memory mechanisms observed during training. B Distributions of learning rates and resulting memory mechanisms as a function of the delay period. Each line represents a distinct learning rate; each run is initialized with a different random seed. C Power-law fits shown on a log-log scale, where the $y$-axis indicates the maximum learnable delay and the $x$-axis the corresponding critical learning rate. Each dot denotes the cutoff delay value beyond which learning fails (see Methods for how these values are estimated). D-F Same as in panels A-C, but for tasks with post-reaction windows. G The phase diagram of mechanisms predicted by the toy model. H-I Phase diagrams corresponding to panels B-E, where color intensity reflects the proportion of trials converging to each of the three dynamical phases. As predicted by the theoretical analysis in G, the original task (H) reveals three distinct regimes, whereas the inclusion of a post-reaction window (I) biases the model toward limit cycle dynamics. To generate these phase diagrams, we trained over 65,000 RNNs across varying learning rates, delay intervals ($T_{\rm delay}$), and random seeds (see Table S1). For full parameter details, refer to Section S4.
@@
In Appendix S2.2, we compute an analytical loss function for the limit cycle model in the limits $T_{\rm resp}=T_{\rm delay}$ and $T_{\rm resp}\ll T_{\rm delay}$. We illustrate the latter case in Fig. S2B, note the exact correspondence between the analytical curve and the numerical simulations. For both cases, the loss function has a global optimum at $r=\frac{1}{2T_{\rm delay}}$. However, it has a kink around the global minimum and quickly flattens to a constant value for $r\leq\frac{1}{2(T_{\rm delay}+T_{\rm resp})}$. Hence, if the learning rate is larger than some critical value, the oscillations around the global minimum can throw the parameter $r$ to a value $r<\frac{1}{2(T_{\rm delay}+T_{\rm resp})}$, terminating the training process due to the zero gradient despite the fact that the model cannot solve the task. Our analysis reveals that this maximum learning rate scales following $\alpha_{\rm LC}^{*}\sim T_{\rm delay}^{-\beta_{\rm LC}}$ where $\beta_{\rm LC}\in[2,3]$. Notably, given that $\beta_{\rm LC}\leq\beta_{\rm SP}$, the limit cycle solution has a better scaling than the slow-point one, providing a theoretical explanation for our observation in Fig. 2. Specifically, since the RNN in Fig. 2 was trained using a learning rate right at the transition (compare to Fig. S1), it is not surprising that a slow point mechanism eventually evolved into a limit cycle. Moreover, this theoretical prediction is also consistent with the experiments in Fig. 3C, where increasing the delay period for a fixed learning rate resulted in more RNNs finding limit cycle based solutions. Hence, both learning rate and the duration of the delay can influence the learned latent mechanism during training.
Further insights from toy models– The distinct scaling of $\beta_{\rm LC}$ and $\beta_{\rm SP}$ also highlights the existence of an interesting phenomenon: a phase space of mechanisms that can be learned by RNNs. This phase space (which will be discussed in Fig. 4G) depends on the delay time (a task parameter) and the learning rate (an optimization parameter). On the other hand, it is important to not over-interpret the toy model results. Specifically, increasing the dimensionality of the latent dynamical system can allow more efficient solutions to the task, and it may be possible for RNNs to utilize other (diverse) mechanisms for storing memories (e.g., those using inherent transient dynamics in high-dimensional neural networks (Jaeger, 2002; Liu et al., 2025; Maass et al., 2002; Ichikawa & Kaneko, 2021)). Fortunately, the toy models we discussed utilize mechanisms that are observed in practice (Figs. 1, 2, and S1), and therefore can offer further insights into learning dynamics in more complicated architectures. Below, we provide one example case, and show empirically that similar phase spaces emerge in full-rank RNNs trained on delayed cue discrimination tasks.
3.4 Empirical scaling analyses of latent mechanisms
As a final step, we tested the theoretical predictions of the toy models, including the proposed scaling laws and the phase space of latent mechanisms, by training full-rank RNNs on delayed cue-discrimination tasks (Figs. 4, S3, S4, and S5). Similar to low-rank RNNs, full-rank RNNs also learned to solve the task using two distinct mechanisms (Fig. 4A; Fig. S3A,B). Solutions involving sustained oscillations exhibited minimal power at low frequencies, as the dynamics were dominated by high-frequency components (Fig. S3C). Based on this observation, we developed a mechanism discrimination index that automatically classified the learned RNNs into three groups: (i) networks relying on slow-point manifolds, (ii) networks employing limit cycles, and (iii) networks that failed to perform the task (see Fig. S4 for representative examples).
Large-scale RNN experiments reproduced the toy model predictions–We then applied this classification to examine whether different mechanisms consistently emerged across varying delay durations. In line with theoretical predictions, increasing the length of the delay period preferentially led to the emergence of limit cycles (Fig. 4B). These oscillatory solutions disappeared beyond a critical learning rate for a given delay, with the transition scaling as $\beta_{\mathrm{LC}}=2.72\pm 0.07$ (Fig. 4C), in close agreement with predictions (see Fig. S3D-E for the procedure used to compute critical delays). For shorter delays, RNNs predominantly developed slow-point manifolds, with the transition scaling as $\beta_{\mathrm{SP}}=4.05\pm 0.10$, also matching the theoretical range (Fig. S3). However, when we trained RNNs on a modified task variant that included a post-reaction window, slow-point solutions were no longer observed, and only oscillatory dynamics emerged (Fig. 4D-F). To summarize these transitions, we plotted phase diagrams (Fig. 4G-I) derived from both the theoretical model and the full-rank RNN experiments, capturing the conditions under which each mechanism emerged and their associated scaling laws.
Control experiments–Finally, we performed two additional control experiments. In Fig. S5, we repeated the analysis using RNNs with slower intrinsic dynamics (i.e., larger $\tau$), and in Fig. S6, we tested a version of the task that no longer required memory maintenance. In the latter case, the dependence on task duration nearly vanished, and the scaling observed in Fig. 4 was substantially weakened ($\beta=0.38\pm 0.02$). These findings reinforce the idea that the computational mechanisms adopted by RNNs are not fixed but shaped by the task’s demands. In particular, memory duration, learning rate, and structural constraints (refer to Q3 above) jointly determine the preferred mechanism.
@@ 4 Discussion
In this work, we set out to answer three fundamental questions about the population-level mechanisms that can support short-term memory in neural networks.
What mechanisms support memory maintenance through sequential neural activity? In the memory tasks studied here, neural activation sequences consistently emerged as core components of the solutions. These sequences arose through two distinct mechanisms: slow-point manifolds and limit cycles. While the use of slow-points to approximate sequences is well aligned with prior intuitions (Khona & Fiete, 2022), the emergence of limit cycles in this context represents a novel prediction of our framework. Although limit cycles have previously been observed in memory tasks (Pals et al., 2024), they typically appear when periodicity is embedded in the task design. In contrast, the delayed activation and delayed cue-discrimination tasks used here had no explicit periodic components. Thus, the appearance of limit cycles in these tasks suggests an alternative, previously underappreciated, mechanism for generating sustained neural sequences in the absence of external rhythmic input or an oscillatory output.
What determines which mechanism is favored under different memory tasks or training conditions? Our analytical results revealed that the selection between slow-point manifolds and limit cycles depended on both the structure of the task and the learning rate used in the optimization. We later confirmed these predictions in large-scale simulations using full-rank RNNs (Table S1). This finding is particularly striking in light of recent interest in dynamical systems approaches in neuroscience, where considerable effort has gone into identifying and manipulating attractors and manifolds that underlie neural activity (Liu et al., 2024; Vinograd et al., 2024). The sharp divergence in dynamical solutions (slow-points versus limit cycles) emerging from minor changes in task design underscores a critical point: the structure of the task can strongly shape the neural dynamics that emerge.
This raises important questions about how we interpret dynamical structures observed in neural recordings. Can we conclude that one type of solution is more fundamental than another simply because it appears in a specific task variant? Our findings highlight the need for caution. They suggest that conclusions about the nature of neural computations must be grounded not only in observed dynamics but also in a careful analysis of how task parameters may constrain or bias the underlying solutions. What may appear to be a core computational motif could, in fact, reflect a contingent outcome of the task design itself.
@@
How do the learned mechanisms vary with the delay duration? A key insight into this question comes from examining how critical learning rates (beyond which learning ceases to be effective) scale with delay length. In simple delayed-response tasks, we found that the critical learning rate allowing successful training decreases as a power-law function of the delay. As the delay increases, smaller and smaller updates are required, ultimately making training impractically slow for long durations. This constraint may underlie the historical difficulty of capturing long-term dependencies in RNNs, as highlighted in early work (Bengio et al., 1994).
The introduction of gated architectures such as LSTMs and GRUs partially addressed this issue by enabling networks to learn adaptive timescales (Hochreiter, 1997). These architectures could perhaps bypass the power-law constraint by dynamically modulating their internal time-scales, and thereby extend their memory retention periods. However, even if successful, this flexibility would lead to trade-offs. Extending timescales indefinitely slows down network responses, limiting their utility for rapid decision-making. In addition, expanding the parameter space to accommodate longer memory spans can complicate optimization and destabilize learning (Dinc et al., 2025b). Unlike biological systems, which operate under biophysical constraints such as membrane time constants and action potentials, artificial networks must learn timescales through training, which is not always possible even if the underlying network is theoretically capable of solving a particular task.
Overall, by framing the challenge through dynamical systems theory, our work highlights a core trade-off between memory duration, learning speed, and network stability. Understanding this trade-off offers a principled explanation for the persistent difficulty of learning long-term dependencies and opens the door to new strategies that balance biological plausibility, computational efficiency, and learning performance.
Concluding remarks—Our findings carry important implications for experimental neuroscience. We have shown that both slow-points and limit cycles can serve as effective mechanisms for generating sequential neural activity. However, in the context of a dynamic brain, where synaptic plasticity continually reshapes functional connectivity (Ebrahimi et al., 2022), it remains unclear how to determine whether observed sequences are supported by slow-points or limit cycles. More broadly, how can we distinguish between dynamical solutions that produce equivalent activity patterns during a trial, but diverge in their behavior once the trial ends? To our knowledge, this question (concerning equivalence of distinct latent mechanisms subserving task execution, but later diverging after the task conclusion) has not been explicitly addressed in experimental designs within systems neuroscience. Tackling it would require a shift in how experiments are conceptualized and conducted. As a starting point, our work illustrates how (even minor) manipulations of task structure can lead to distinct underlying mechanisms.
Beyond neuroscience, the phenomenon of staged learning dynamics and algorithmic phase transitions has been widely documented in artificial neural networks, including Transformers trained on modular arithmetic (Liu et al., 2022; Nanda et al., 2023), language modeling (Wei et al., 2022; Lubana et al., 2024), and in-context learning (Raventós et al., 2024). Within this growing literature, our study of RNNs provides a complementary perspective, grounded in dynamical systems theory, that connects computational behavior to interpretable latent mechanisms. Recent efforts aim to formalize learning progress in feedforward networks using quantities analogous to order parameters in physics (Park et al., 2024). Our results go a step further by identifying literal phase transitions between qualitatively distinct computational strategies in dynamical systems. Future work may explore how these mechanisms can be implemented or approximated by feedforward architectures.